OpenStudy (anonymous):

evaluate the area of an ellipse 4x^2 +y^2 +4 by using double integration

OpenStudy (turingtest):

typo? 4x^2 +y^2 =4 perhaps?

OpenStudy (anonymous):

i don't understand

OpenStudy (turingtest):

4x^2+y^2+4 is not an ellipse it's not even a graph can you see that?

OpenStudy (anonymous):

it was given in a test and now i have to do it and submit tmr

OpenStudy (turingtest):

they must have meant 4x^2+y^2=4 right? that would be a graph of a ellipse what you typed is just an expression because it has no equals sign, do you not agree?

OpenStudy (turingtest):

well, assuming it is\[4x^2+y^2=4\]I would look at the first quadrant:so it looks like polar coordinates are a nice choice, in which case we have\[D=\left\{ (r,\theta)|0\le r\le1+\cos\theta,0\le\theta\le\frac{\pi}{2} \right\}\]as I said, this is a quarter if the ellipse, so we need to multiply our integral by 4 to get the whole thing\[4\int_{0}^{\frac{\pi}{2}}\int_{0}^{1+\cos\theta}dA=4\int_{0}^{\frac{\pi}{2}}\int_{0}^{1+\cos\theta}rdrd\theta\]to integrate you may need the identity\[\cos^2\theta=\frac{1}{2}(1+\cos(2\theta))\]